OpenStudy (anonymous):

solve the equation M=19r^2h/11 for r in terms of M and h. assume r=M and h are all positive

OpenStudy (lgbasallote):

\[M = \frac{19r^2 h}{11}\] first, multiply both sides by 11 \[\implies 11M = 19r^2 h\] now divide both sides by 19 \[\implies \frac{11M}{19} = r^2 h\] now divide both sides by h \[\implies \frac{11M}{19h} = r^2\] finally.. take the square root to get r does that help?

OpenStudy (anonymous):

yes, thanks

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

help in this one solve 3z+3t=b-1 for z

OpenStudy (lgbasallote):

first subtract both sides by 3t \[\implies 3x = b - 1 - 3t\] now divide both sides by 3

OpenStudy (anonymous):

so it will be z=b-1-t